# 22footnotetext: Corresponding author

### v1

@@ Abstract.
Mamba, a special case of the State Space Model, is gaining popularity as an alternative to template-based deep learning approaches in medical image analysis. While transformers are powerful architectures, they have drawbacks, including quadratic computational complexity and an inability to efficiently address long-range dependencies. This limitation affects the analysis of large and complex datasets in medical imaging, where there are many spatial and temporal relationships. In contrast, Mamba offers benefits that make it well-suited for medical image analysis. It has linear time complexity, which is a significant improvement over transformers. In sequence modeling tasks, computational complexity grows linearly with the length of the input sequence. Mamba processes longer sequences without attention mechanisms, enabling faster inference and requiring less memory. Mamba also demonstrates strong performance in merging multimodal data, improving diagnosis accuracy and patient outcomes. The paper’s organization allows readers to appreciate Mamba’s capabilities in medical imaging step by step. We begin with clear definitions of relevant concepts regarding SSMs and concept models, including S4, S5, and S6. We then explore Mamba architectures, including pure Mamba, U-Net variants, and hybrid models that combine Mamba with convolutional networks, transformers, and Graph Neural Networks. Subsequent sections cover Mamba optimizations, techniques such as weakly supervised and self-supervised learning, scanning mechanisms, and a detailed analysis of applications across various tasks. We provide an overview of available datasets and several experimental results regarding Mamba’s efficacy in different domains. Furthermore, we detail the challenges and limitations of Mamba, along with other interesting aspects and possible future directions. The final subsection explains the importance of Mamba in medical imaging and provides an analysis and conclusions regarding its usage and enhancement measures. This review aims to demonstrate the transformative potential of Mamba in overcoming existing barriers within medical imaging while paving the way for innovative advancements in the field. A comprehensive list of Mamba architectures applied in medical field, reviewed in this work is available on Github [^4] https://github.com/Madhavaprasath23/Awesome-Mamba-Papers-On-Medical-Domain.
copyright: nonecopyright: acmlicensedjournalyear: 2024publicationmonth: 10
1. Introduction
@@
In recent few decades, there has been a remarkable improvement in the field of medicine through the application of machine learning (Shehab et al., 2022) as well as deep learning (Suzuki, 2017). The initial architectures of neural networks like Convolutional Neural Networks (CNNs) (Li et al., 2021) played a pivotal role in better image segmentation (Kayalibay et al., 2017), classification (Li et al., 2014; Raghaw et al., 2024a), and object detection (Li et al., 2019). Medical images are complex, but CNN’s were able to analyze 3D structures in a 2D plane and so proved useful in biomedical image computing especially into image segmentation (Ronneberger et al., 2015), tumor detection (Choudhury et al., 2020), organ segmentation (Zhao et al., 2019), and disease diagnosis imaging (Chen et al., 2019). CNNs have been applied extensively to medical imaging tasks, namely segmentation, classification and reconstruction. One weakness, however, is that they can lack when sequencing data, or multitasking which requires long-range dependencies. For example, in the area of medical image segmentation, the use of CNNs may not perform well as one would expect since they may not be able to model super-resolution inter-dependence of an image and its parts.
@@
Some of the drawbacks of CNNs have been addressed by transformers (Vaswani et al., 2017; Raghaw et al., 2024b) as advancements of technologies that have better sequential data processing and long-range dependencies. Still, there are some disadvantages as well. The main problem is the scaling of computed attention which grows quadratically with the sequence length, thus makes the use of such attention costly and hard on sequences that are very long. Moreover, many additional resources and data are usually needed, which is quite a problem if one has to work in a resource-restricted environment such as in the medical domain. In regard to the shortcomings of classical CNNs and transformers, there has been noticeable progress in research on different types of models which could efficiently represent the long sequences and their intricate dependencies. Of late, State Space Models (SSMs) (Gu et al., 2021a) have gained much attention as one of such alternatives as the Mamba (Gu and Dao, 2024) model.
@@
Figure 1. Evolution of Mamba from State Space Models (SSMs)
@@
Mamba, aims to address the problems related to modern deep learning techniques. Selective state spaces are employed to quickly assimilate vast lengths of sequences, combine various modes and command extensive yet practical resolutions. The architecture of Mamba incorporates selective scan mechanism and hardware aware algorithm which support high efficiency in storage and computation of the intermediate results. This helps Mamba perform very well in some tasks such as medical image segmentation (Xing et al., 2024; Wu et al., 2024a; Wang et al., 2024a), classification (Gong et al., 2024; Yue and Li, 2024; Nasiri-Sarvi et al., 2024), synthesis, registration and reconstruction (Huang et al., 2024b; Zheng and Zhang, 2024; Li et al., 2024b) where long range dependency and high complexity is involved. Mamba has performed promisingly well in the biomedical field, especially in the fields of biomedical imaging, genomics and processing clinical notes. Thus, the model comes in handy in capturing long range and multi-modal data oriented tasks which involve subtle relationships and dependencies between units of information. Figure 1 shows the timeline of Mamba’s evolution over time, starting from HiPPO (Gu et al., 2020) and SSMs such as Linear State Space Layer (LSSL) (Gu et al., 2021b), S4 (Gu et al., 2021a), Diagonal State Space (DSS) (Gupta et al., 2022) , S4D (Gu et al., 2022), S5 (Smith et al., 2023), S4ND (Nguyen et al., 2022), Hungry Hungry Hippos (H3) (Fu et al., 2023) to Mamba (Gu and Dao, 2024). It also includes variants of Mamba that were created as the model evolved.
@@
Figure 2. Distribution of Research Papers Utilizing Mamba in Medical Domain
The pie chart, as depicted in Figure 2 illustrates the distribution of research papers utilizing the Mamba framework across various tasks in the medical domain. The chart is divided into five segments, each representing a specific task and its corresponding percentage contribution to the total number of papers. Moreover, Figure 3 illustrates the fluctuation in the number of publications related to Mamba in medical domain over the period from December 2023 to September 2024. A notable surge in research activity is evident in March and April 2024.
There exist several survey papers on Mamba. However, these might do either of the following – works (Patro and Agneeswaran, 2024; Qu et al., 2024) cover the framework broadly or are restricted to its use only in the vision domain (Xu et al., 2024; Liu et al., 2024c; Zhang et al., 2024e). It is worth noting that only (Heidari et al., 2024) provided a review of Mamba which was focused on its uses in medical domain. However, our survey paper is more extensive and detailed than those of (Heidari et al., 2024). In particular, the present work emphasizes the analysis of public resources such as medical datasets, and presents some empirical data on the applicability of Mamba within medical practice, including various resources and interventions to be utilized within Mamba in a healthcare context. In addition to this, we include the latest research and developments of Mamba architectures for medical image analysis. Moreover, we have structured our work in such a way that readers will appreciate the organizational spans of Mamba including its strengths, weaknesses and prospects within the medical arena.
Figure 3. Publication Trend of Mamba Research in the Medical Domain
2. Core Concepts of SSM
In the realm of deep learning, Transformers have consistently dominated in both Computer Vision (CV) and Natural Language Processing (NLP) tasks. The self-attention (Vaswani et al., 2017) mechanism within Transformers has greatly improved the understanding of these modalities by generating an attention matrix from the query, key, and value vectors. While the attention matrix is beneficial, it suffers from quadratic time complexity. The recent advancements, such as FlashAttention by Dao et al. (Dao et al., 2022; Dao, 2023) and linear attention (Katharopoulos et al., 2020a), have addressed this issue by reducing the time complexity. For instance, in linear attention, the key is multiplied by the value instead of the query, and the softmax function is replaced with a similarity function. Mamba developed by Gu et al. (Gu and Dao, 2024) further mitigates this problem by transforming the quadratic time complexity into linear time complexity in a recurrent manner. Mamba is the first model without attention to match the performance of a very strong Transformer. The core concepts of Mamba and its derivation from SSM are explained in the following sections.
2.1. State Space Models
State Space Models (SSMs) uses an approach similar to Kalman filter (Kalman, 1960). SSMs convert a one-dimensional input sequence $u(t)$ into an N-dimensional continuous latent state $x(t)$, which is then projected into a 1D output signal $y(t)$. The entire process of a state space model can be represented as shown in Equation 1 and Equation 2:
(1) $x^{\prime}(t)=\textbf{{A}}x(t)+\textbf{{B}}u(t)$
(2) $y(t)=\textbf{{C}}x(t)+\textbf{{D}}u(t)$

### v3

# A Comprehensive Survey of Mamba Architectures for Medical Image Analysis: Classification, Segmentation, Restoration and Beyond

@@ Abstract
Mamba, a special case of the State Space Model, is gaining popularity as an alternative to template-based deep learning approaches in medical image analysis. While transformers are powerful architectures, they have drawbacks, including quadratic computational complexity and the inability to efficiently address long-range dependencies. This limitation affects the analysis of large and complex datasets in medical imaging, where there are many spatial and temporal relationships. In contrast, Mamba offers benefits that make it well suited for medical image analysis. It has linear time complexity, which is a significant improvement over transformers. In sequence modeling tasks, computational complexity grows linearly with the length of the input sequence. Mamba processes longer sequences without attention mechanisms, enabling faster inference and requiring less memory. Mamba also demonstrates strong performance in merging multimodal data, improving diagnosis accuracy and patient outcomes. The paper’s organization allows readers to appreciate Mamba’s capabilities in medical imaging step by step. We begin with clear definitions of relevant concepts regarding SSMs and concept models, including S4, S5, and S6. We then explore Mamba architectures, including pure Mamba, U-Net variants, and hybrid models that combine Mamba with convolutional networks, transformers, and Graph Neural Networks. Subsequent sections cover Mamba optimizations, techniques such as weakly supervised and self-supervised learning, scanning mechanisms, and a detailed analysis of applications across various tasks. We provide an overview of available datasets and several experimental results regarding Mamba’s efficacy in different domains. Furthermore, we detail the challenges and limitations of Mamba, along with other interesting aspects and possible future directions. The final subsection explains the importance of Mamba in medical imaging and provides an analysis and conclusions regarding its usage and enhancement measures. This review aims to demonstrate the transformative potential of Mamba in overcoming existing barriers within medical imaging while paving the way for innovative advancements in the field. A comprehensive list of Mamba architectures applied in the medical field, reviewed in this work, is available on Github [^4] https://github.com/Madhavaprasath23/Awesome-Mamba-Papers-On-Medical-Domain.
keywords:
Convolutional Neural Networks; Transformers; State Space Models; Mamba; Medical Image Analysis
journal: Computer Methods and Programs in Biomedicine
1 Introduction
@@
In recent few decades, there has been a remarkable improvement in the field of medicine through the application of machine learning [1] as well as deep learning [2]. The initial architectures of neural networks like Convolutional Neural Networks (CNNs) [3] played a pivotal role in better image segmentation [4], classification [5, 6], and object detection [7]. Medical images are complex, but CNNs were able to analyze 3D structures in a 2D plane and so proved useful in biomedical image computing especially in image segmentation [8], tumor detection [9], organ segmentation [10], and disease diagnosis imaging [11]. CNNs have been applied extensively to medical imaging tasks, namely segmentation, classification, and reconstruction. One weakness, however, is that they can lack when sequencing data, or multitasking, which requires long-range dependencies. For example, in the area of medical image segmentation, the use of CNNs may not perform well as one would expect since they may not be able to model super-resolution inter-dependence of an image and its parts.
@@
Some of the drawbacks of CNNs have been addressed by transformers [12, 13] as advancements of technologies that have better sequential data processing and long-range dependencies. Still, there are some disadvantages as well. The main problem is the scaling of computed attention which grows quadratically with the sequence length, thus makes the use of such attention costly and hard on sequences that are very long. Moreover, many additional resources and data are usually needed, which is quite a problem if one has to work in a resource-restricted environment such as in the medical domain. In regard to the shortcomings of classical CNNs and transformers, there has been noticeable progress in research on different types of models which could efficiently represent the long sequences and their intricate dependencies. Of late, State Space Models (SSMs) [14] have gained much attention as one of such alternatives as the Mamba [15] model.
@@
Figure 1: Evolution of Mamba from State Space Models (SSMs)
@@
Mamba, aims to address the problems related to modern deep learning techniques. Selective state spaces are employed to quickly assimilate vast lengths of sequences, combine various modes, and command extensive yet practical resolutions. The architecture of Mamba incorporates selective scan mechanism and hardware aware algorithm which support high efficiency in storage and computation of the intermediate results. This helps Mamba perform very well in some tasks such as medical image segmentation [16, 17, 18, 19], classification [20, 21, 22], registration and reconstruction [23, 24, 25] where long range dependency and high complexity is involved. Mamba has performed promisingly well in the biomedical field, especially in the fields of biomedical imaging, genomics and processing clinical notes. Thus, the model comes in handy in capturing long range and multi-modal data oriented tasks which involve subtle relationships and dependencies between units of information. Figure 1 shows the timeline of Mamba’s evolution over time, starting from HiPPO [26] and SSMs such as Linear State Space Layer (LSSL) [27], S4 [14], Diagonal State Space (DSS) [28] , S4D [29], S5 [30], S4ND [31], Hungry Hungry Hippos (H3) [32] to Mamba [15]. It also includes variants of Mamba that were created as the model evolved.
The pie chart, as depicted in Figure 2 illustrates the distribution of research papers utilizing the Mamba framework across various tasks in the medical domain. The chart is divided into five segments, each representing a specific task and its corresponding percentage contribution to the total number of papers. Moreover, Figure 3 illustrates the fluctuation in the number of publications related to Mamba in the medical domain during the period from December 2023 to September 2024. A notable surge in research activity is evident in March and April 2024.
4.1%8.2%12.2%22.4%53.1%Image RegistrationImage RestorationImage Classification MiscellaneousImage Segmentation
Figure 2: Distribution of Research Papers Utilizing Mamba in Medical Domain
Existing surveys in this domain either cover the framework broadly [33, 34] or are restricted to its use only in the vision domain [35, 36, 37]. It is worth noting that only [38] provided a review of Mamba which was focused on its uses in the medical domain. However, our survey paper is more extensive and detailed than that of [38]. In particular, the present work emphasizes the analysis of public resources, such as medical datasets and presents some empirical data on the applicability of Mamba within medical practice, including various resources and interventions to be utilized within Mamba in a healthcare context. Beyond this, we incorporate recent advances in self-supervised and multimodal learning within Mamba architectures, demonstrating their potential to improve medical image analysis beyond traditional supervised approaches. We also address practical challenges related to deploying Mamba models in real-world clinical settings, including hardware limitations and generalization across diverse medical datasets. In addition to this, we include the latest research and developments of Mamba architectures for medical image analysis. Furthermore, we explore future directions aimed at improving computational efficiency, adapting scanning mechanisms for non-causal visual data, and expanding Mamba’s usability in real-time and edge computing scenarios. Moreover, we have structured our work in such a way that readers will appreciate the organizational spans of Mamba including its strengths, weaknesses, and prospects within the medical arena. Figure 4 displays the taxonomy of this survey, providing a comprehensive overview of the topics covered.
Figure 3: Publication Trend of Mamba Research in the Medical Domain
A Comprehensive Survey of Mamba Architectures for Medical Image Analysis: Classification, Segmentation, Restoration and Beyond Conclusion Future directions Key findings Significance for the field Discussion Emerging Areas:
Mamba 2[39], xLSTM[40]
Limitations Datasets BraTS2023[41, 42, 43], AIIB2023[44, 45], CRC-500[16], ISIC2017[46], ISIC2018[47, 48], CVC-ClinicDB[49], 3D Abdomen CT[50], 2D Abdomen MRI[51], Endoscopy[52], Microscopy[53], ACDC MRI Cardiac[54], Synapse Multi-Organ Abdominal CT55footnotemark: 5, LiTS[55], Montgomery and Shenzhen[56], Brain MRI Multiple Sclerosis[57], ADNI[58, 59], Otoscopy[60], PathMNIST[61, 62], CAMELYON16[63, 64], Colorectal Cancer Histopathology[65], SynthRAD[66], FastMRI[67], Low-Dose CT Image and Projection[68] Applications in Various Medical Domains Miscellaneous:
Prompt-Mamba[69], ClinicalMamba[70], P-Mamba[71], MD-Dose[72], Vivim[73], VM-DDPM[74]
Medical Image Registration:
MambaMorph[66], VMambaMorph[75]
Medical Image Restoration/Reconstruction:
FDVM-Net[24], MambaMIR[23], MambaDFuse[25], FusionMamba[76]
Medical Image Classification:
nnMamba[20], MedMamba[21], Vim4path[22], Microscopic-Mamba[77]
Medical Image Segmentation:
SegMamba[16], H-Vmunet[17], LKM-UNet[18], Mamba-UNet[78], Weak Mamba UNet[79], LightM-UNet[80], UltraLight-VM UNet[81], T-Mamba[82], HC-Mamba[83], Semi-Mamba-UNet[84], ViM-UNet[85], U-Mamba[86], Mamba-HUNet[87], $U$-Mamba[88], CAMS-Net[89], MUCM-Net[90]
Techniques and Adaptations Multimodal Learning: Fusion Mamba[76], MGI[91], GFE-Mamba[92] Self Supervised Learning: Vim4Path[22], MGI[91], MambaMIM[93], CMViM[94] Semi-Supervised Learning: Semi-Mamba-UNet[84] Weakly Supervised Learning: Weak-Mamba-UNet[79] Mamba Optimizations Lightweight and Efficient:
LightM UNet[80], UltraLight VM-UNet[81], MUCM-Net[90], LightCF-Net[95], MiM-ISTD[96]
Scanning BiDirectional Scan [97], Selective Scan 2D [98], Spatiotemporal Selective Scan [73], Zigzag Scan [99], Local Scan [100], Efficient 2D Scan [101], Continuous 2D Scan [102], Three Directional Scan [16], Pixelwise and Patchwise Scan [18], Omnidirectional Selective Scan [103], Hierarchical Scan [104], Multi-Head Scan [105], Multi-Path Scan [106], 3D BiDirectional Scan [107], Bidirectional Slice Scan [108] Mamba architectures Hybrid Architectures:
Mamba with convolution: U-Mamba[86], Mamba-U-Net[78], HC-Mamba[83], TS-Mamba[109], nnMamba[20], Swin-UMamba[110]
Mamba with Attention and Transformers: Weak-Mamba-UNet[79], TP-Mamba[111], HMT-UNet[112]
Mamba with Recurrence: VMRNN[113]
Miscellaneous: Vim4Path[22], MambaMIL[114]
Variants of U-Net:
VM-UNet[115], ViM-UNet[85], VM-UNet-V2[116], LKM-UNet[18], H-vmunet[17]
Pure Mamba:
ViM[97], VMamba[98], Plain Mamba[102], EVSS[101], MambaND[117], MHS-VM[105], Local Mamba[100], Weak-Mamba-UNet[79], TMamba[82], H-vmunet[17], UltraLight VM-UNet[81], Light M-UNet[80], SegMamba[16], LKM-UNet[18]
Core Concepts of SSM Selective Structured State Space Models (S6) [15] Simplified State Space Layers for Sequence Modeling (S5) [30] Structured State Space Sequence Models (S4) [14] State Space Models
Figure 4: Structural Taxonomy of Survey Content
2 Core Concepts of SSM
In the realm of deep learning, Transformers have consistently dominated in both Computer Vision (CV) and Natural Language Processing (NLP) tasks. The self-attention [12] mechanism within Transformers has greatly improved the understanding of these modalities by generating an attention matrix from the query, key, and value vectors. While the attention matrix is beneficial, it suffers from quadratic time complexity. The recent advancements, such as FlashAttention by Dao et al. [118, 119] and linear attention [120], have addressed this issue by reducing the time complexity. For instance, in linear attention, the key is multiplied by the value instead of the query, and the softmax function is replaced with a similarity function. Mamba developed by Gu et al. [15] further mitigates this problem by transforming the quadratic time complexity into linear time complexity in a recurrent manner. Mamba is the first model without attention to match the performance of a very strong Transformer. The core concepts of Mamba and its derivation from SSM are explained in the following sections.
2.1 State Space Models
State Space Models (SSMs) uses an approach similar to Kalman filter [121]. SSMs convert a one-dimensional input sequence $u(t)$ into an N-dimensional continuous latent state $x(t)$, which is then projected into a 1D output signal $y(t)$. The entire process of a state space model can be represented as shown in Equation 1 and Equation 2:
$x^{\prime}(t)=\textbf{{A}}x(t)+\textbf{{B}}u(t)$ (1)
$y(t)=\textbf{{C}}x(t)+\textbf{{D}}u(t)$ (2)
@@ 2.1 State Space Models
2.2 Structured State Space Sequence Models (S4)
S4 proposed by Gu et al. [14] demonstrates how to efficiently compute all forms of the SSM: the recurrent representation, and the convolutional representation. Additionally, S4 employs a bilinear method for discretizing the parameters of the SSM, converting the state space parameter A into an approximation $\overline{\textbf{{A}}}$. In S4, the parameter $D$ from the original state space model is either set to 0 or used as a residual connection. The discrete SSM can be expressed in its recurrent form as shown in Equation 3 and Equation 4. The recurrent representation of the equation below can be scanned in parallel using a prefix sum operation, as the current input does not explicitly depend on the previous hidden layer. This makes parallel scan achievable, thereby reducing time complexity from linear to logarithmic.
$x_{k}=\overline{\textbf{{A}}}x_{k-1}+\overline{\textbf{{B}}}u_{k}\;\;\;\;\;\overline{\textbf{{A}}}=(\mathbf{I}-\Delta/2\cdot\textbf{{A}})^{-1}(\mathbf{I}+\Delta/2\cdot\textbf{{A}})$ (3)
$y_{k}=\overline{\textbf{{C}}}x_{k}\;\;\;\;\;\;\;\overline{\textbf{{B}}}=(\mathbf{I}-\Delta/2\cdot\textbf{{{A}}})^{-1}(\Delta\textbf{{B}})\;\;\;\;\;\;\;\overline{\textbf{{C}}}=\textbf{{C}}$ (4)
$x_{0}=\overline{\textbf{{B}}}u_{0}\;\;\;\;\;\;x_{1}=\overline{\textbf{{AB}}}u_{0}+\overline{\textbf{{B}}}u_{1}\;\;\;\;\;\;x_{2}=\overline{\textbf{{A}}}^{2}\overline{\textbf{{B}}}u_{0}+\overline{\textbf{{AB}}}u_{1}+\overline{\textbf{{B}}}u_{2}$ (5)
$y_{0}=\overline{\textbf{{CB}}}u_{0}\;\;\;\;\;\;y_{1}=\overline{\textbf{{CAB}}}u_{0}+\overline{\textbf{{CB}}}u_{1}\;\;\;\;\;\;y_{2}=\overline{\textbf{{CA}}}^{2}\overline{\textbf{{B}}}u_{0}+\overline{\textbf{{CAB}}}u_{1}+\overline{\textbf{{CB}}}u_{2}$ (6)
$y_{k}=\overline{\textbf{{CA}}}^{k}\overline{\textbf{{B}}}u_{0}+\overline{\textbf{{CA}}}^{k-1}\overline{\textbf{{B}}}u_{1}+......+\overline{\textbf{{CAB}}}u_{k-1}+\overline{\textbf{{CB}}}u_{k}$ (7)
$y=\overline{\textbf{{K}}}*u.$ (8)
$\overline{\textbf{{K}}}\;\in\;\mathbb{R}^{L}:=\kappa_{L}(\overline{\textbf{{A}}},\overline{\textbf{{B}}},\overline{\textbf{{C}}}):=(\overline{\textbf{{CA}}}^{i}\overline{\textbf{{B}}})_{i\in[L]}=(\overline{\textbf{{CB}}},\overline{\textbf{{CAB}}},...,\overline{\textbf{{CA}}}^{L-1}\overline{\textbf{{B}}}).$ (9)
$\textbf{(HiPPO Matrix)}\;\;\;\;\;\;\;\;\;\;\textbf{{A}}_{nk}=-\left\{\begin{matrix}(2n+1)^{1/2}(2k+1)^{1/2}\;\;if\;n>k\\ n+1\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;if\;n=k\\ 0\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;if\;n<k\end{matrix}\right$. (10)
S4 addresses the limitations of transformers by implementing these strategies. This empowers SSMs to excel in tasks requiring long-range dependencies such as Path-X [122] . In contrast, Transformers [123, 124] in Path-X exhibit accuracy below 50% (worse than random guessing), whereas S4 achieves approximately 80% accuracy.
2.3 Simplified State Space Layers for Sequence Modeling (S5)
S5 proposed by Smith et al. [30], extends from S4 with similar initialization conditions, but enhances the architecture by adopting a Multiple Input Multiple Output (MIMO) approach. Notably, S5 introduces a learnable time scale parameter ($\Delta$), replacing the fixed parameter used in S4. Parameters in S5 are discretized using the Zero Order Hold (ZOH) method as mentioned in Equation 11, providing a refined parameter system compared to S4.
$\overline{\mathbf{\Lambda}}=e^{\mathbf{\Lambda}\Delta},\;\;\;\;\;\overline{\textbf{{B}}}=\mathbf{\Lambda}^{-1}(\overline{\mathbf{\Lambda}}-\mathbf{I})\tilde{\textbf{{B}}},\;\;\;\;\;\overline{\textbf{{C}}}=\tilde{\textbf{{C}}},\;\;\;\;\;\overline{\textbf{{D}}}=\tilde{\textbf{{D}}}$. (11)
In terms of computation, S5 employs a fully recurrent connection with parallel scanning , as detailed in Equation 3 and Equation 4. The authors highlight that a smaller latent space employs SSM to do parallel scanning where an associative operation is used in between during offline settings. This characteristic positions S5 for both online and offline processing, highlighting its utility in recurrent tasks within the time domain.
2.4 Selective Structured State Space Models (S6)
S6 introduced by Gu et al. [15] initializes the parameter A from S4D which employs a diagonalized matrix structure for A. Extended from S5, S6 incorporates SSMs within the Mamba Architecture. S6 builds upon the foundational assumptions of previous SSMs by leveraging a projection of the input function (using a linear layer) for initializing parameters $B$ and $C$. Notably, S6 also applies this projection to the step size parameter, i.e., $\Delta$. To enhance computational efficiency, S6 implements faster recurrence connections using operations on the Static Random Access Memory (SRAM) of the GPU, with storage on the High Bandwidth Memory (HBM) similar to the principles outlined in FlashAttention mechanism by Dao et al. [118] and FlashAttention-2 mechanism by Dao et al. [119].
Algorithm 1 SSM (S4)
$\textbf{Input :}\;\;x\;:\;(B,L,D)$
$\textbf{Output :}\;\;y\;:\;(B,L,D)$